Vector products and groups of bundle automorphisms

Wilson Sutherland, 1st November, 1996

Informal abstract: I’ll begin by recalling the relation between vector products and almost complex structures on spheres, then describe a similar thing — a special case of a simple question that comes up in connection with gauge groups: when is a vector bundle ξ equivalent to its tensor product ξ ⊗ λ with a line bundle λ? This seems to be not quite as simple as it looks, but if someone can see that it really is simple, the talk may collapse at an early stage!